Please Proove tge question (1+x) ^-n=1-nx, where x

Grade:11

1 Answers

Aditya Gupta
5 years ago
by the generalised binomial theorem,
(1+x) ^-n=1-nx+n(n+1)/2!*x^2 – …..
now assuming x is very very less than 1 so that the second and higher orders of x are negligible, we write
(1+x) ^-n=1 – nx, provided x is small. QED